# Observing structural disorder induced interacting topological phase in an atom array

### v2

@@ Acknowledgements.
In the presence of structural disorder, it is clear to see that for a typical sample $C=\{x_{j}:j=1,2,\dots,2N\}$, the Hamiltonian no longer preserves the inversion symmetry, i.e., $U_{R}H^{\textrm{S}}(C)U_{R}^{-1}\neq H^{\textrm{S}}(C)$. As a result, $P_{\mathrm{S}}$, as defined in Eq. (S3), is not restricted to quantization at 00 or $0.5$ any more, and can take other values. However, if we consider an ensemble of all possible lattice configurations, an average inversion symmetry remains. Specifically, the Hamiltonian ensemble is defined as $\mathcal{E}_{H}\equiv\left\{H^{\textrm{S}}(C):C\in\mathcal{E}_{C}\right\}$, where $\mathcal{E}_{C}$ is an ensemble containing all possible random lattice configurations. In our case, any random lattice configuration $C$ and its inversion partner $\mathcal{R}C$ emerge in this ensemble with equal probability. As a result, the Hamiltonian $H(C)$ and its inversion conjugate partner $U_{R}H^{\textrm{S}}(C)U_{R}^{-1}=H^{\textrm{S}}(\mathcal{R}C)$ appear with the same probability. This Hamiltonian ensemble thus preserves an average inversion symmetry 32, 33, 34. With this observation, we identify the topological property of the Hamiltonian $H^{\textrm{S}}(C)$ and $H^{\textrm{S}}(\mathcal{R}C)$ as a whole, defining the polarization as
@@
where $D(S_{0})$ and $U(S_{0})$ are the corresponding matrices for the lattice configuration $S_{0}$ with coordinates of $x_{i}(S_{0})$. Given that $H^{\textrm{S}}(C)|u_{i}\rangle=E_{i}|u_{i}\rangle$, we have $H^{\textrm{S}}(\mathcal{R}C)U_{R}|u_{i}\rangle=E_{i}U_{R}|u_{i}\rangle$, indicating that $U_{R}|u_{i}\rangle$ is the corresponding eigenstate of $H^{\textrm{S}}(\mathcal{R}C)$, or $U(\mathcal{R}C)=U_{R}U(C)$. Based on this, we derive that
@@
[…]
indicating that $P_{\mathrm{S}}(C)$ can only take the values of either 00 or $0.5$.
@@
Let $|u_{\mathrm{L}}(C)\rangle$ and $|u_{\mathrm{R}}(C)\rangle$ be the two edge states of the single-particle Hamiltonian $H^{\mathrm{S}}(C)$ for a random configuration $C$ under open boundary conditions. They mainly reside at the left and right edges, respectively, and their corresponding eigenenergies are $E_{\mathrm{L}}(C)$ and $E_{\mathrm{R}}(C)$. Due to the breaking of the inversion symmetry, $E_{\mathrm{L}}(C)\neq E_{\mathrm{R}}(C)$ for each realized configuration. For the inversion conjugate Hamiltonian $H^{\mathrm{S}}(\mathcal{R}C)$, $|u_{\mathrm{R}}(\mathcal{R}C)\rangle=U_{R}|u_{\mathrm{L}}(C)\rangle$, and $|u_{\mathrm{L}}(\mathcal{R}C)\rangle=U_{R}|u_{\mathrm{R}}(C)\rangle$ are its eigenstates corresponding to energy $E_{\mathrm{L}}(C)$ and $E_{\mathrm{R}}(C)$, respectively. They are mainly localized at the right and left edges, respectively. Consider the local DOS, $\rho(i=1,E)(C)$ and $\rho(i=2N,E)(C)$ [$\rho(i=1,E)(\mathcal{R}C)$ and $\rho(i=2N,E)(\mathcal{R}C)$], at two edges at the energy $E$ for the configuration $C$ (the inversion partner $\mathcal{R}C$), for simplicity, we only include contributions from the edge state: At the left edge,
@@
and similarly for the local DOS at the right edge. For the two configurations, the average of the two local DOS at the left edge is $[\rho(i=1,E)(C)+\rho(i=1,E)(\mathcal{R}C)]/2$, which is clearly equal to the one at the right edge, $[\rho(i=2N,E)(C)+\rho(i=2N,E)(\mathcal{R}C)]/2$. Thus, it can be concluded that the two edge states are statistically degenerate. For clarity, we plot the numerically computed configuration averaged local DOS in Fig. S1B, illustrating that it preserves the inversion symmetry, and two red peaks at the left and right edges have the same energy.
@@
In the single-particle case, van der Waals interactions contribute an onsite energy of $-V_{i}^{\text{vdW}}$ at site $i$ ($i=1,2,\dots,2N$). In a regular lattice with periodic boundary conditions, this onsite potential is independent of the site index. Consequently, it does not alter the eigenstates of the Hamiltonian, leaving the topology unaffected. However, under open boundary conditions as in our experiment, the potentials at edges differ from in the bulk due to the absence of one neighbor for the edge sites, as shown in Fig. S1(C). Despite this difference, the total Hamiltonian still respects the inversion symmetry, ensuring that two edge states, if they exist, are degenerate. In the figure, we see that in the bulk $V_{i}^{\text{vdW}}/h\approx 0.7$ MHz, which is comparable to the intracell hopping energy of $0.77$ MHz in the regular case. However, at the edges, the potential is only $0.02$ MHz. This means if there were edge states in the regular lattice, we would have observed a peak of occupancy at two edges near zero detuning by microwave spectroscopy. However, our experiments (see Fig. 2(B) in the main text) do not show such a peak, implicating the absence of edge modes.
In the disordered case, for a typical configuration, $V_{i}^{\text{vdW}}$ is dependent on the site index even under periodic boundary conditions. However, we have proven in the previous subsection that the topology of the ensemble of Hamiltonians in disordered lattices can still be characterized by the generalized polarization $P_{\textrm{S}}(C)$, and edge states, if they exist, are statistically degenerate. Figure S1(D) displays the $V_{i}^{\text{vdW}}$ averaged over the $15$ random configurations used in experiments. We see that similarly to the regular case, $V_{i}^{\text{vdW}}$ at edge sites is about $0.04$ MHz. The local DOS in Figs. S1A and S1B show that the peaks at the edges appear at $-0.07$ MHz, whose absolute value is larger than the averaged onsite energies at the edges, possibly due to non-vanishing occupancy probabilities in the bulk sites for edge states. The existence of edge states is confirmed by our experiments, illustrating a peak of occupancy at two edges near zero detuning (see Fig. 2(B) in the main text).
@@
In this subsection, we will focus on the case at or near half-filling ($N$, $N-1$, or $N+1$ atoms in the Rydberg $p$ state). It is more convenient to write the Hamiltonian in Eq. (1) in the main text as a spin model
@@
The Hamiltonian $H_{\mathrm{XXZ}}$ respects two spin-rotation symmetries along $x$ and $y$ respectively represented by $\hat{R}_{x}=\prod_{j=1}^{2N}e^{-i\pi\sigma_{j}^{x}/2}$ and $\hat{R}_{y}=\prod_{j=1}^{2N}e^{-i\pi\sigma_{j}^{y}/2}$, the time-reversal symmetry represented by $\hat{T}=\prod_{j=1}^{2N}\sigma_{j}^{y}\kappa$ with $\kappa$ the complex conjugate operator, and a symmetry represented by $\hat{\mathcal{S}}=\prod_{j=1}^{2N}\sigma_{j}^{x}\kappa$. The model $\hat{H}$ preserves the total spin along $z$, i.e., $[\hat{H},\sigma^{z}]=0$ with $\sigma^{z}=\sum_{i}\sigma_{i}^{z}$, which corresponds to the $U(1)$ symmetry.
@@
We first demonstrate that the hard-core boson model in Eq. (1) in the main text or the spin model in Eq. (S12) can be mapped to a Hamiltonian of interacting fermions 26, 36 through the inverse Jordan-Wigner transformation, i.e.,
@@
To characterize the topology of the spin model Eq. (12), we define a $\mathbb{Z}_{2}$ topological invariant 49, 50
$P_{\mathrm{M}}=\left[\frac{1}{2\pi}\mathrm{Im}\ln\langle\Psi|\hat{\mathcal{P}}_{\mathrm{M}}|\Psi\rangle\right]\text{ mod }1,$ (S18)
in analogy to the case of single particles, where $|\Psi\rangle$ is the ground state of the spin model in the symmetry sector with $\sigma^{z}=0$ under periodic boundary conditions and $\hat{\mathcal{P}}_{\mathrm{M}}=\prod_{j=1}^{2N}e^{-\frac{\pi i}{N{d}}x_{j}\sigma_{j}^{z}}$ is the twist operator with $x_{j}$ denoting the position of the $\lfloor(j+1)/2\rfloor$th unit cell. It has been proven that any one of the above symmetries for $H_{\mathrm{XXZ}}$ ensures that $P_{\mathrm{M}}$ can only take the values of either 00 or $0.5$ 36, making it a well-defined topological invariant. For a topologically trivial phase, $P_{\mathrm{M}}=0$, while for a nontrivial phase, $P_{\mathrm{M}}=0.5$. Analogous to the single-particle case, van der Waals interactions introduce the $H_{\mathrm{Z}}$ term, which breaks all the above protecting symmetries. In the following, we will prove that either an inversion symmetry or an average inversion symmetry can protect the quantization of the topological invariant for the ground state in the symmetry sector with $\sigma^{z}=0$.
@@
For bosonic SPT phases, inversion symmetry plays the role of time-reversal symmetry based on the crystalline equivalence principle 66, 67, leading to a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ classification in 1D 68, provided the system also respects the $U(1)$ symmetry. Here we will present a proof demonstrating that inversion symmetry can protect the quantization of $P_{\mathrm{M}}$ in the spin model. Note that an alternative proof has been provided for a spin-1 model in Ref. 50.
Specifically, suppose that the spin model has an inversion symmetry, i.e., $\mathcal{U}_{R}\hat{H}\mathcal{U}_{R}^{-1}=\hat{H}$, where $\mathcal{U}_{R}$ realizes the inversion operation or $\mathcal{U}_{R}\sigma_{j}^{\nu}\mathcal{U}_{R}^{-1}=\sigma_{2N+1-j}^{\nu}$ ($\nu=x,y,z$). We note when the $H_{\mathrm{Z}}$ term is sufficiently large, $|\Psi\rangle$ as the ground state of this model in the symmetry sector with $\sigma^{z}=0$ is no longer assured to be the global ground state of the total Hamiltonian $\hat{H}$. In the regular case as well as for most of the random configurations, we find that this state remains the ground state of the total Hamiltonian. In fact, ground states are prepared in a fixed symmetry sector in our experiments and we have not encountered any serious trouble with them not being that of the total Hamiltonian. In addition, we assume that $|\Psi\rangle$ is not degenerate so that $\mathcal{U}_{R}|\Psi\rangle=\lambda|\Psi\rangle$ with $\lambda=1$ or $-1$. We now derive that
@@
where we have used the fact that $\sigma^{z}|\Psi\rangle=0$. This indicates that $\langle\Psi|\hat{\mathcal{P}}_{\mathrm{M}}|\Psi\rangle$ is real so that $P_{\mathrm{M}}$ has to be quantized to 00 or $0.5$.
@@
Similar to the single-particle case, in the presence of structural disorder, for a typical configuration $C$, the inversion symmetry is broken, i.e., $\mathcal{U}_{R}\hat{H}(C)\mathcal{U}_{R}^{-1}\neq\hat{H}(C)$. However, the Hamiltonian ensemble $\mathcal{E}_{\hat{H}}\equiv\left\{\hat{H}(C):C\in\mathcal{E}_{C}\right\}$ possesses an average inversion symmetry. In other words, the Hamiltonian $\hat{H}(C)$ and its inversion conjugate partner $\mathcal{U}_{R}\hat{H}(C)\mathcal{U}_{R}^{-1}=\hat{H}(\mathcal{R}C)$ appear with the same probability. With this average symmetry, we define a $\mathbb{Z}_{2}$ topological invariant in Eq. (2) as in the main text. We now prove that it is enforced to be quantized by the average symmetry for the ground state $|\Psi(C)\rangle$ of the configuration $C$ and the ground state $|\Psi(\mathcal{R}C)\rangle$ of the configuration $\mathcal{R}C$ in the symmetry sector with $\sigma^{z}=0$. Based on the property that $\sigma^{z}|\Psi(C)\rangle=0$, we show that
@@
Thus, $\sum_{S\in\{C,\mathcal{R}C\}}\langle\Psi_{S}|\hat{\mathcal{P}}_{\mathrm{M}}(S)|\Psi_{S}\rangle$ is real, and $P_{\textrm{M}}(C)$ is quantized to 00 or $0.5$.
In the above proof, we have made use of the fact that $\sigma^{z}$ is conserved corresponding to the $U(1)$ symmetry. As such, the symmetry class of the system is given by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$, where the spatial inversion symmetry is treated as an antiunitary $\mathbb{Z}_{2}$ internal symmetry through the crystalline equivalence principle 66, 67. Based on previous classification results 35, the average SPT phase protected by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$ in 1D is given by
@@
respectively. Here $N_{C}$ denotes the number of elements in the ensemble $\mathcal{E}_{C}$. To show that $\overline{n}_{1}=\overline{n}_{2N}$, we only need to prove that they are equal when $\mathcal{E}_{C}=\{C,\mathcal{R}C\}$, thanks to the average inversion symmetry. The following derivation constitutes the proof
@@
III.2.6 S3.2.6. Effects of van der Waals interactions
@@
In the previous subsections, we have established the concept of the average inversion SPT phase by showing that the $\mathbb{Z}_{2}$ topological invariant is enforced to be quantized, and the ground states in the half-filling subspace is two-fold statistically degenerate.

### v3

@@ Acknowledgements.
In the presence of structural disorder, it is clear to see that for a typical sample $C=\{x_{j}:j=1,2,\dots,2N\}$, the Hamiltonian no longer preserves the inversion symmetry, i.e., $U_{R}H^{\textrm{S}}(C)U_{R}^{-1}\neq H^{\textrm{S}}(C)$. As a result, $P_{\mathrm{S}}$, as defined in Eq. (S4), is not quantized at 0 or $0.5$ any more and can take other values. However, if we consider the ensemble of all possible lattice configurations, an average inversion symmetry remains. Specifically, the Hamiltonian ensemble is defined as $\mathcal{E}_{H}\equiv\left\{H^{\textrm{S}}(C):C\in\mathcal{E}_{C}\right\}$, where $\mathcal{E}_{C}$ is an ensemble containing all possible random lattice configurations. In our case, any random lattice configuration $C$ and its inversion partner $\mathcal{R}C$ emerge in this ensemble with equal probability. As a result, the Hamiltonian $H(C)$ and its inversion conjugate partner $U_{R}H^{\textrm{S}}(C)U_{R}^{-1}=H^{\textrm{S}}(\mathcal{R}C)$ appear with the same probability. This Hamiltonian ensemble thus preserves an average inversion symmetry 13, 14, 16. We thus can identify the topological property of the Hamiltonian $H^{\textrm{S}}(C)$ and $H^{\textrm{S}}(\mathcal{R}C)$ as a whole, defining the polarization as
@@
where $D(S_{0})$ and $U(S_{0})$ are the corresponding matrices for the lattice configuration $S_{0}$ with coordinates of $x_{i}(S_{0})$. Given that $H^{\textrm{S}}(C)|u_{i}\rangle=E_{i}|u_{i}\rangle$, we have $H^{\textrm{S}}(\mathcal{R}C)U_{R}|u_{i}\rangle=E_{i}U_{R}|u_{i}\rangle$, indicating that $U_{R}|u_{i}\rangle$ is the corresponding eigenstate of $H^{\textrm{S}}(\mathcal{R}C)$, or $U(\mathcal{R}C)=U_{R}U(C)$. Thus we find
@@
$P_{\mathrm{S}}(C)=\left\{\frac{1}{2\pi}\mathrm{Im}\ln\left[e^{-i2\pi\frac{1}{2N{d}}\sum_{i=1}^{2N}x_{i}(C)}\det\left(U(C)^{\dagger}D(C)U(C)\right)+\textrm{c.c.}\right]\right\}\textrm{ mod }1,$ (S10)
indicating that $P_{\mathrm{S}}(C)$ can only take the values of either 0 or $0.5$.
@@
Let $|u_{\mathrm{L}}(C)\rangle$ and $|u_{\mathrm{R}}(C)\rangle$ be the two edge states of the single-particle Hamiltonian $H^{\mathrm{S}}(C)$ for a random configuration $C$ under open boundary conditions. They reside mainly at the left and right edges, respectively, and their corresponding eigenenergies are $E_{\mathrm{L}}(C)$ and $E_{\mathrm{R}}(C)$. Due to the breaking of the inversion symmetry, $E_{\mathrm{L}}(C)\neq E_{\mathrm{R}}(C)$ for each realized configuration. For the inversion conjugate Hamiltonian $H^{\mathrm{S}}(\mathcal{R}C)$, $|u_{\mathrm{R}}(\mathcal{R}C)\rangle=U_{R}|u_{\mathrm{L}}(C)\rangle$, and $|u_{\mathrm{L}}(\mathcal{R}C)\rangle=U_{R}|u_{\mathrm{R}}(C)\rangle$ are its eigenstates corresponding to energy $E_{\mathrm{L}}(C)$ and $E_{\mathrm{R}}(C)$, respectively. They are mainly localized at the right and left edges, respectively. Consider the local DOS, $\rho(i=1,E)(C)$ and $\rho(i=2N,E)(C)$ $\rho(i=1,E)(\mathcal{R}C)$ and $\rho(i=2N,E)(\mathcal{R}C)$], at two edges at the energy $E$ for the configuration $C$ (the inversion partner $\mathcal{R}C$), for simplicity, we only include contributions from the edge state: At the left edge,
@@
and similarly for the local DOS at the right edge. For the two configurations, the average of the two local DOS at the left edge is $\rho(i=1,E)(C)+\rho(i=1,E)(\mathcal{R}C)]/2$, which is clearly equal to the one at the right edge, $\rho(i=2N,E)(C)+\rho(i=2N,E)(\mathcal{R}C)]/2$. Thus, it can be concluded that the two edge states are statistically degenerate. For better clarity, we plot the numerically computed configuration averaged local DOS in Fig. S2b, illustrating that it preserves the inversion symmetry, and the two red peaks at the left and right edges have the same energy.
@@
In the single-particle case, van der Waals interactions contribute an onsite energy of $V_{i}^{\text{vdW}}$ at site $i$ ($i=1,2,\dots,2N$). In a regular lattice with periodic boundary conditions, this onsite potential is independent of the site index. Consequently, it does not alter the eigenstates of the Hamiltonian, leaving the topology unaffected. However, under open boundary conditions as in our experiment, the potentials at edges differ from in the bulk due to the absence of one nearest neighbor for the edge sites, as shown in Fig. S2c. Despite this difference, the total Hamiltonian still respects the inversion symmetry, ensuring that two edge states, if they exist, are degenerate. In the figure, we see that in the bulk $V_{i}^{\text{vdW}}/h\approx 0.7$ MHz, which is comparable to the intracell hopping energy of $0.77$ MHz in the regular case. However, at the edges, the potential is only $0.02$ MHz. This means if there were edge states in the regular lattice, we would have observed a peak of occupancy at two edges near zero detuning by microwave spectroscopy. However, our experiments (see Fig. 2b in the main text) do not show such a peak, implicating the absence of edge modes.
In the disordered case, for a typical configuration, $V_{i}^{\text{vdW}}$ depends on the site index even under periodic boundary conditions. However, we have shown in the previous subsection that the topology of the ensemble of Hamiltonians in disordered lattices can still be characterized by the generalized polarization $P_{\textrm{S}}(C)$, and edge states, if they exist, are statistically degenerate. Figure S2d displays the $V_{i}^{\text{vdW}}$ averaged over the $15$ random configurations used in our experiments. We see that similarly to the regular case, $V_{i}^{\text{vdW}}$ at edge sites is about $0.04$ MHz. The local DOS in Figs. S2a and S2b show that the peaks at the edges appear at 0.07-0.07 MHz, whose absolute value is larger than the averaged onsite energies at the edges, possibly due to non-vanishing occupancy probabilities in the bulk sites for edge states. The existence of edge states is confirmed by our experiments, illustrating a peak of occupancy at two edges near zero detuning (see Fig. 2b in the main text).
@@
In this subsection, we will focus on the case at or near half-filling ($N$, N1N-1, or $N+1$ atoms in the Rydberg $p$ state). In this case it is more convenient to write the Hamiltonian in Eq. (1) in the main text as a spin model
@@
The Hamiltonian $H_{\mathrm{XXZ}}$ respects two spin-rotation symmetries along $x$ and $y$ respectively represented by $\hat{R}_{x}=\prod_{j=1}^{2N}e^{-i\pi\sigma_{j}^{x}/2}$ and $\hat{R}_{y}=\prod_{j=1}^{2N}e^{-i\pi\sigma_{j}^{y}/2}$, the time-reversal symmetry represented by $\hat{T}=\prod_{j=1}^{2N}\sigma_{j}^{y}\kappa$ with $\kappa$ the complex conjugate operator, and a symmetry represented by $\hat{\mathcal{S}}=\prod_{j=1}^{2N}\sigma_{j}^{x}\kappa$. The model $\hat{H}$ preserves the total spin along $z$, i.e., $\hat{H},\sigma^{z}]=0$ with $\sigma^{z}=\sum_{i}\sigma_{i}^{z}$, which corresponds to the $U(1)$ symmetry.
@@
We first demonstrate that the hard-core boson model in Eq. (1) in the main text or the spin model in Eq. (S13) can be mapped to a Hamiltonian of interacting fermions 32, 39 through the inverse Jordan-Wigner transformation, i.e.,
@@
To characterize the topology of the spin model Eq. (12), we define a $\mathbb{Z}_{2}$ topological invariant 62, 63
[…]
in analogy to the case of single particles, where $|\Psi\rangle$ is the ground state of the spin model in the symmetry sector with $\sigma^{z}=0$ under periodic boundary conditions and $\hat{\mathcal{P}}_{\mathrm{M}}=\prod_{j=1}^{2N}e^{-\frac{\pi i}{N{d}}x_{j}\sigma_{j}^{z}}$ is the twist operator with $x_{j}$ denoting the position of the $\lfloor(j+1)/2\rfloorth$ unit cell. It has been proven that any one of the above symmetries for $H_{\mathrm{XXZ}}$ ensures that $P_{\mathrm{M}}$ can only take the values of either 0 or $0.5$ 39, making it a well-defined topological invariant. For a topologically trivial phase, $P_{\mathrm{M}}=0$, while for a nontrivial phase, $P_{\mathrm{M}}=0.5$. Analogous to the single-particle case, van der Waals interactions introduce the $H_{\mathrm{Z}}$ term, which breaks all the above protecting symmetries. In the following, we will prove that either an inversion symmetry or an average inversion symmetry can protect the quantization of the topological invariant for the ground state in the symmetry sector with $\sigma^{z}=0$.
@@
For bosonic SPT phases, inversion symmetry plays the role of time-reversal symmetry based on the crystalline equivalence principle 79, 80, leading to a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ classification in 1D 8, provided the system also respects the $U(1)$ symmetry. Here we will present a proof demonstrating that inversion symmetry can protect the quantization of $P_{\mathrm{M}}$ in the spin model. Note that an alternative proof has been provided for a spin-1 model in Ref. 63.
Specifically, suppose that the spin model has an inversion symmetry, i.e., $\mathcal{U}_{R}\hat{H}\mathcal{U}_{R}^{-1}=\hat{H}$, where $\mathcal{U}_{R}$ realizes the inversion operation or $\mathcal{U}_{R}\sigma_{j}^{\nu}\mathcal{U}_{R}^{-1}=\sigma_{2N+1-j}^{\nu}$ ($\nu=x,y,z$). We note when the $H_{\mathrm{Z}}$ term is sufficiently large, $|\Psi\rangle$ as the ground state of this model in the symmetry sector with $\sigma^{z}=0$ is no longer assured to be the global ground state of the total Hamiltonian $\hat{H}$. In the regular case as well as for most of the random configurations, we find that this state remains the ground state of the total Hamiltonian. In fact, ground states are prepared in a fixed symmetry sector in our experiments and we have not encountered any serious trouble with them not being that of the total Hamiltonian. In addition, we assume that $|\Psi\rangle$ is not degenerate so that $\mathcal{U}_{R}|\Psi\rangle=\lambda|\Psi\rangle$ with $\lambda=1$ or 1-1. We now derive that
@@
where we have used the fact that $\sigma^{z}|\Psi\rangle=0$. This indicates that $\langle\Psi|\hat{\mathcal{P}}_{\mathrm{M}}|\Psi\rangle$ is real so that $P_{\mathrm{M}}$ has to be quantized to 0 or $0.5$.
@@
[…]
@@
Thus, $\sum_{S\in\{C,\mathcal{R}C\}}\langle\Psi_{S}|\hat{\mathcal{P}}_{\mathrm{M}}(S)|\Psi_{S}\rangle$ is real, and $P_{\textrm{M}}(C)$ is quantized to 0 or $0.5$.
In the above proof, we have made use of the fact that $\sigma^{z}$ is conserved corresponding to the $U(1)$ symmetry. As such, the symmetry class of the system is given by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$, where the spatial inversion symmetry is treated as an antiunitary $\mathbb{Z}_{2}$ internal symmetry through the crystalline equivalence principle 79, 80. Based on previous classification results 16, the average SPT phase protected by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$ in 1D is given by
@@
respectively. Here $N_{C}$ denotes the number of elements in the ensemble $\mathcal{E}_{C}$. To show that $\overline{n}_{1}=\overline{n}_{2N}$, we only need to prove that they are almost equal when $\mathcal{E}_{C}=\{C,\mathcal{R}C\}$, thanks to the average inversion symmetry. The following derivation constitutes the proof
@@
III.2.6 S3.2.6. Effects of van der Waals interactions on ground-state degeneracy
@@
In the previous parts, we have established the concept of the average inversion SPT phase by showing that the $\mathbb{Z}_{2}$ topological invariant is enforced to be quantized, and the ground states in the half-filling subspace are two-fold statistically degenerate.